# Improving the mean-field approximation in continuous models of population dynamics with nonlocal dispersal: applications to vegetation pattern formation

### v1

@@ Abstract
Spot patterns, in which vegetation patches form a hexagonal lattice, are frequent in nature and could serve as an early-warning indicator of abrupt vegetation collapses. Consequently, they have been intensively studied using both individual-based models and density-based field equations. Yet, the relationship between these two approaches remains unclear, particularly in scenarios where vegetation dynamics exhibit strong long-range spatial correlations and traditional mean-field approximations fail. To solve this issue, we develop a new method that refines mean-field approximations by describing both the dynamics of the biomass density field and its spatial correlations. This new approach harnesses the strengths of both individual and density-based mdoels, treating spatial correlations explicitly and allowing for the identification of spatial instabilities resulting in periodic patterns. Our results indicate that this new approximation predicts the parameter regimes where regular periodic patterns emerge more accurately than mean-field models, suggesting that it could provide a more robust framework to perform further nonlinear analysis.
@@ 1 Introduction
Plants in water-limited ecosystems provide a good system to investigate regular pattern formation in populations of sessile and semi-sessile organisms [10]. Several studies using aerial and satellite images from different locations have identified a variety of patterns in these systems, including gaps, stripes, rings, labyrinths, and spots [11, 12, 13, 14]. These patterns are primarily shaped by environmental factors such as gradients in resource availability, interactions between plants and other organisms, and local topography [15, 16, 17, 18, 14, 19]. For example, along a gradient of increasing aridity, vegetation pattern shapes shift from gaps to labyrinths and ultimately to spots [12]. Therefore, these spot patterns, where approximately circular patches of vegetation form a hexagonal lattice in a matrix of bare soil [20], could indicate proximity to non-vegetated states [8]. This potential role as an early-warning indicator of desertification processes makes spot patterns ecologically relevant. However, the long timescales governing vegetation dynamics and the limited availability of long-term datasets pose challenges for empirically investigating the ecological significance of vegetation patterns [but see 21, 22]. Alternatively, researchers have used spatially explicit theoretical models to explain both how spot patterns could form and the types of desertification processes they might precede [23, 24, 25, 26, 27]. Broadly, existing models can be classified into two categories: discrete individual-based (IBM), and continuous density-based models.
IBMs describe a population as a spatial distribution of point-like organisms that reproduce and die according to probabilistic rules that encode different interactions and dispersal mechanisms [28, 29]. These models are built at the level of individual plants, which makes it possible to study spatial patterns resolving the patch structure, and are easy to simulate computationally. However, the computational overhead associated with these simulations scales rapidly with the system size and the complexity of interactions, and analytical techniques to study them are limited to providing predictions of population size and its spatial correlations at equilibrium [30, 31, 32]. In continuous models, on the other hand, the vegetation biomass is a field that changes in space and time following a differential equation [33, 34]. This approach is easier to treat mathematically, which makes it possible to predict the dynamics of certain spatial processes, such as the direction of ‘desertification’ fronts [35]. However, density-based models do not resolve individual plants, which limits how much they inform about patch composition. This limitation turns them inappropriate for studying random and irregular aggregated patterns [36] unless they consider environmental heterogeneity [37, 38, 39, 40] or demographic fluctuations explicitly [41, 42, 43, 44].
@@ 2.1 Description of the stochastic dynamics
To keep our analysis as simple as possible, we consider a simple logistic-like birth-death dynamics in continuous time [36]. The $i$-th individual, located at $\mathbf{x}_{i}$, reproduces and disperses a new organism at location ${\mathbf{x}}$ at a rate
@@
where the first term on the right side corresponds to baseline mortality, and the second term accounts for density-dependent effects due to intraspecific competition. $\phi_{\mathrm{C}}$ is the competition kernel that weighs the influence of an organism at ${\mathbf{x}}_{j}$ on the focal individual at ${\mathbf{x}}_{i}$. The shape of the competition kernel is a key ingredient in determining the type of spatial patterns that can form [46, 51]. In the simulations, we will use a Gaussian with mean zero and standard deviation $r_{c}$ and a top-hat competition kernel of the form
@@
In both cases, $r_{c}$ defines the characteristic spatial scale of the competition. We performed all IBM simulations in a square domain of lateral length $L=20$, using periodic boundary conditions and demographic rates $d/b=0.1$, with $b=3.05$ and $g=1$.
@@ 2.2 Spatial moment dynamics equations
The discrete dynamics in Section 2.1 can be approximated at the macroscale in terms of the moments of the population spatial pattern. For the kind of spatial point process considered in the IBM, the first spatial moment, $Z_{1}(t)$, is the average density of individuals. The second spatial moment, $Z_{2}(\bm{\xi},t)$, is defined as the average density of pairs of individuals separated by a displacement $\bm{\xi}$. Similarly, the third spatial moment, $Z_{3}(\bm{\xi},\bm{\xi}^{\prime},t)$, corresponds to the average density of triplet of individuals separated by displacements $\bm{\xi}$ and $\bm{\xi}^{\prime}$, respectively [31, 52, 53]. For a given spatial pattern $p(\mathbf{x},t)$, the spatial moments are defined as
$Z_{1}(t)$ $=\frac{1}{A}\int_{\Omega}p(\mathbf{x},t)\mathrm{d}\mathbf{x}$ (2.4a)
$Z_{2}(\bm{\xi},t)$ $=\frac{1}{A}\int_{\Omega}p(\mathbf{x},t)\left(p(\mathbf{x}+\bm{\xi},t)-\delta(\bm{\xi})\right)\mathrm{d}\mathbf{x}$ (2.4b)
$Z_{3}(\bm{\xi},\bm{\xi}^{\prime},t)$ $=\frac{1}{A}\int_{\Omega}p(\mathbf{x},t)\left(p(\mathbf{x}+\bm{\xi},t)-\delta(\bm{\xi})\right)\times$ (2.4c)
@@
in which the delta funtions in Eq. (2.4b) and Eq. (2.4c) prevent counting pairs and triplets with repeated organisms. Notice that we can obtain the radial correlation function from the definition of the second spatial moment (2.4b) as
@@ 2.3 Field equation for the biomass density field
where we have omitted the time dependence on the right side. Notice that, similarly to what we obtained in the spatial moment dynamics equations in Section 2.2, Eq. (2.19) does not provide a closed description for the dynamics of the population density because it also includes its spatial covariance. We could write an equation for the dynamics of this spatial covariance and perform a moment closure approximation, as we did with the spatial moment dynamics equations. This approach is now much more complicated because the plant density is a field rather than a scalar quantity obtained from a spatial average. Therefore, we explore two alternative ways of treating these spatial correlations.
@@ 2.3.2 Beyond mean-field: introducing the effect of spatial covariances
All the quantities on the right side of Eq. (2.28) can be computed from IBM simulations. Alternatively, these quantities can also be obtained using SMD equations, which is computationally more efficient and makes it possible to have a closed dynamical equation for the vegetation density field without performing IBM simulations. Using the definitions of the spatial moments, we can rewrite (2.28) as,
@@ 3.1 Spatial patterns in the discrete model
Figure 1: Spatial moment dynamics results using a Gaussian competition kernel. A) Radial correlation function at displacement zero, $C(0)$ [36]. B) Total biomass density normalized by the mean-field expectation, $n_{\mathrm{MF}}=(b-d)/g$, as a function of the dispersal and competition ranges, $s$ and $r_{c}$ respectively. C) Radial correlation function for three combinations of $s$ and $r_{c}$ representative of each type of spatial pattern. Each curve is labeled with a symbol that indicates the values of $s$ and $r_{c}$ in (B). $b$, $d$, and $g$ with $d/b=0.1$ and $g=1$
@@
Previous work has identified four possible types of spatial patterns depending on how these two scales compare to each other [36]. When both ranges are large, the dynamics are effectively non-spatial, and the IBM converges to the classical logistic model. However, when the dispersal scale decreases, while keeping competition long-ranged, individuals tend to clump, which leads to smaller population sizes than in the non-spatial case. Likewise, when the competition range decreases while keeping dispersal long-range, individuals tend to overdisperse, which increases the total population size. Finally, when both competition and dispersal are short-range, newborns cannot escape from competition with their ancestors and the population goes extinct. We recover these results integrating the SMD approximation in Eqs. (2.7)-(2.9) with the moment closure in Eq. (2.10) (Fig. 1a, b). In none of the scenarios discussed above, however, the spatial patterns are periodic (see radial correlation functions in Fig. 1c). Therefore, any of the continuous approximations we derived for the plant biomass density will return a homogeneous solution.
@@
To investigate whether regular patterns exist in our model, we solved the SMD equations using a top-hat kernel as defined in Eq. (2.3) and compared these predictions with results from simulations of the discrete IBM. The IBM simulations show the emergence of regular patterns when the dispersal range is short and the competition range long (Fig. 2). This result is in good agreement with previous work using various nonlocal extensions of the Fisher KPP equation with top-hat competition kernels [46, 51]. We find that the SMD equations accurately predict the value of the radial correlation function at pairwise distance zero (top panels in Fig. 3b, c), which is a proxy for local crowding. The SMD equations, however, predict less accurately the population density. Specifically, when the competition range is large, and the dispersal range is short, the SMD equations underestimate the total population size resulting from the IBM simulations (central panels in Fig. 3a, b).
@@
Figure 2: A) Long-time periodic pattern snapshot with dispersal range $s=0.15$ and competition range $r_{c}=2.2$ (inside the region of high aggregation in the $C(0)$ heatmaps in Fig. 3). B) Fourier transform of the pattern snapshot.
@@
This discrepancy in the total population density comes from the inaccurate prediction of the full radial correlation function provided by the SMD equations. In particular, SMD does not fully capture the amplitude of the oscillations in the radial correlation functions measured from IBM simulations (rightmost panel in the bottom row of Fig. 3b, c). This amplitude, especially when the radial correlation function becomes negative, is critical to determine how larger the population density grows when regular patterns form. More negative values of $C(r)$ indicate that the regions between clusters are less populated, which favors the growth of more individuals within each aggregate [47].
Note that the moment closure method employed here to approximate the third moment in terms of the first two moments inherently relies on two key assumptions. First, the third moment must be positive, as it represents triplet density and is calculated from the product of three non-negative densities (Eq. (2.4c)). Therefore, the closure approximation must also remain non-negative. The second assumption is the dynamical invariance under relabeling, implying that relabeling identical individuals should not alter the system dynamics [55]. Different closure methods balance these assumptions, with the power-2 closure (used in our computations) offering an optimal trade-off between them. However, in certain parameter regimes, these assumptions may not hold, leading to notable discrepancies between IBM and SMD [59, 55] as we observed in the radial correlation function of the periodic patterns (diamond panels in Fig. 3). Additionally, truncating the hierarchy of spatial moments at the second order may limit model accuracy for spatial structures with significant amounts of information in spatial moments at higher orders.
Figure 3: Discrete model analysis using a top-hat competition kernel. A) Snapshots of spatial patterns. From left to right: segregated, uniform, and regular aggregated. Notice that to facilitate the visualization of the pattern spatial structure, the aggregated and uniform patterns are shown only in a small region of the spatial domain. B) From top to bottom, SMD predictions for the radial correlation function at zero displacement, $C(0)$, the total density of plants normalized by the homogenous expectation, and the radial correlation function for the three combinations of $s$ and $r_{c}$ used in (A) and indicated by the coordinates of the red symbols in the $C(0)$ panel. and carrying capacity derived from spatial moment dynamics. C) Same properties as B) but computed from direct simulations of the stochastic IBM.

### v2

@@ Abstract
Spot patterns, in which vegetation patches form a hexagonal lattice, are frequent in nature and could serve as an early-warning indicator of abrupt vegetation collapses. Consequently, they have been intensively studied using both individual-based models and density-based field equations. Yet, the relationship between these two approaches remains unclear, particularly in scenarios where vegetation dynamics exhibit strong long-range spatial correlations and traditional mean-field approximations fail. To solve this issue, we develop a new method that refines mean-field approximations by describing both the dynamics of the biomass density field and its spatial correlations. This new approach harnesses the strengths of both individual and density-based models, treating spatial correlations explicitly and allowing for the identification of spatial instabilities resulting in periodic patterns. Our results indicate that this new approximation predicts the parameter regimes where regular periodic patterns emerge more accurately than mean-field models, suggesting that it could provide a more robust framework to perform further nonlinear analysis.
@@ 1 Introduction
Plants in water-limited ecosystems provide a good system to investigate regular pattern formation in populations of sessile and semi-sessile organisms [10]. Several studies using aerial and satellite images from different locations have identified a variety of patterns in these systems, including gaps, stripes, rings, labyrinths, and spots [11, 12, 13, 14]. These patterns are primarily shaped by environmental factors such as gradients in resource availability, interactions between plants and other organisms, and local topography [15, 16, 17, 18, 14, 19]. For example, along a gradient of increasing aridity, pattern shapes shift from gaps to labyrinths and ultimately to spots [12]. Therefore, these spot patterns, where approximately circular patches of vegetation form a hexagonal lattice in a matrix of bare soil [20], could indicate proximity to non-vegetated states [8].
This potential role as an early-warning indicator of desertification processes makes spot patterns ecologically relevant. However, the long timescales governing vegetation dynamics and the limited availability of long-term datasets pose challenges for empirically investigating the ecological significance of vegetation patterns [but see 21, 22]. Alternatively, researchers have used spatially explicit theoretical models to explain both how spot patterns could form and the types of desertification processes they might precede [23, 24, 25, 26, 27]. Broadly, existing models can be classified into two categories: discrete individual-based (IBM), and continuous density-based models.
IBMs describe a population as a spatial distribution of point-like organisms that reproduce and die according to probabilistic rules that encode different interactions and dispersal mechanisms [28, 29]. These models are built at the level of individual plants, which makes it possible to study spatial patterns resolving the patch structure, and are easy to simulate computationally. However, the computational overhead associated with these simulations scales rapidly with the system size and the complexity of interactions, and analytical techniques to study them are limited to providing predictions of population size and its spatial correlations at equilibrium [30, 31, 32].
In continuous models, on the other hand, the vegetation biomass is a field that changes in space and time following a differential equation [33, 34]. This approach is easier to treat mathematically, which makes it possible to predict the dynamics of certain spatial processes, such as the direction of ‘desertification’ fronts [35]. However, density-based models do not resolve individual plants, which limits how much they inform about patch composition. This limitation turns them inappropriate for studying random and irregular aggregated patterns [36] unless they consider environmental heterogeneity [37, 38, 39, 40] or demographic fluctuations explicitly [41, 42, 43, 44].
@@ 2.1 Description of the stochastic dynamics
To keep our analysis as simple as possible, we consider a simple birth-death dynamics in continuous time [36]. The $i$-th individual, located at $\mathbf{x}_{i}$, reproduces and disperses a new organism at location ${\mathbf{x}}$ at a rate
@@
where the first term on the right side corresponds to baseline mortality, and the second term accounts for density-dependent effects due to intraspecific competition. $\phi_{\mathrm{C}}$ is the competition kernel that weighs the influence of an organism at ${\mathbf{x}}_{j}$ on the focal individual at ${\mathbf{x}}_{i}$ [51]. The competition kernel models the negative impact of neighbors on the fitness of a focal individual due to, for example, intraspecific competition for shared limited resources. The shape of the competition kernel is a key ingredient in determining the type of spatial patterns that can form [46, 52]. In the simulations, we used two different types of competition kernels. The first type is a Gaussian kernel with mean zero and standard deviation, $r_{c}$. The second is a top-hat competition kernel, defined by the form:
@@
In both cases, $r_{c}$ defines the characteristic spatial scale of the competition.
Note that, while we present a conceptual model and use dimensionless parameters, every model parameter quantifies a process driving plant population dynamics in natural scenarios and can thus be linked to different plant traits. The variance of the dispersal kernel $s$ is a proxy for the distance over which a plant disperses its seeds. The characteristic spatial range of competition, $r_{c}$, defines the distance over which plants’ roots compete for belowground shared resources such as water and minerals. Finally, the intrinsic birth and death rates, $b$ and $d$, represent the rates at which an individual plant reproduces and dies, respectively, while in isolation. Therefore, even though our model does not model a specific plant species, approximated parameter values can be extracted from empirical measurements of dispersal kernels, root systems, or plant life-history studies [53, 54].
The combined effect of birth and death terms in Eq. (2.1) and Eq. (2.2) produces a simple logistic-like birth-death dynamics. We performed all IBM simulations in a square domain of lateral length $L=20$, using periodic boundary conditions. We fixed this domain size to ensure that it is sufficiently large compared to the interaction scales $s$ and $r_{c}$ to observe meaningful patterns within the domain. A summary of key parameters of the model is provided in Table 1.
Symbol Parameter Value
$b$ Intrinsic birth rate 3.05
$d$ Intrinsic death rate 0.305
$s$ Dispersal range 0 – 1.25
$r_{c}$ Spatial range of competition 0 – 4
$g$ Strength of competition 1
Table 1: Model parameters and typical values.
@@ 2.2 Spatial moment dynamics equations
The discrete dynamics in Section 2.1 can be approximated at the macroscale in terms of the moments of the population spatial pattern. For the kind of spatial point process considered in the IBM, the first spatial moment, $Z_{1}(t)$, represents the average plant density within the population [31]. More formally, the first spatial moment for a given vegetation spatial pattern, $p(\mathbf{x},t)$, over a two-dimensional landscape $\Omega$, with area $|\Omega|=A$ is defined as,
$Z_{1}(t)=\frac{1}{A}\int_{\Omega}p(\mathbf{x},t)\mathrm{d}\mathbf{x}.$ (2.4)
The second spatial moment, $Z_{2}(\bm{\xi},t)$, is the average density of pairs of plants separated by a displacement $\bm{\xi}$ [55], which is formally defined as,
$Z_{2}(\bm{\xi},t)=\frac{1}{A}\int_{\Omega}p(\mathbf{x},t)\left(p(\mathbf{x}+\bm{\xi},t)-\delta(\bm{\xi})\right)\mathrm{d}\mathbf{x}.$ (2.5)
Similarly, the third spatial moment, $Z_{3}(\bm{\xi},\bm{\xi}^{\prime},t)$, corresponds to the average density of triplet of plants separated by displacements $\bm{\xi}$ and $\bm{\xi}^{\prime}$, respectively [56]. Again, for a given vegetation spatial pattern $p(\mathbf{x},t)$, the third spatial moment is defined as,
$Z_{3}(\bm{\xi},\bm{\xi}^{\prime},t)$ $=\frac{1}{A}\int_{\Omega}p(\mathbf{x},t)\left(p(\mathbf{x}+\bm{\xi},t)-\delta(\bm{\xi})\right)\times$ (2.6)
@@
where the delta functions in Eq. (2.5) and Eq. (2.6) prevent counting pairs and triplets with repeated plants. Finally, we can obtain the radial correlation function from the definition of the second spatial moment (2.5) as
@@
Note that normalizing the correlation function in Eq. (2.7) by $Z_{1}^{2}(t)$–the density of pairs of plants under complete spatial randomness (alternatively termed uniform patterns)–ensures that $C(r,t)=1$ in the absence of spatial patterns (Fig. 1A, circle panel) [51]. As such, when $C(r,t)>1$, the measured spatial structure has more pairs separated by a distance $r$ than a population with complete spatial randomness. We refer to these types of spatial patterns as clustered spatial patterns (Fig. 1A: Triangle panel). On the other hand, when $C(r,t)<1$, there are fewer plant pairs with a separation distance $r$ than if they were in a spatially random configuration. We refer to these patterns as over-dispersed or segregated spatial patterns (Fig. 1A: Diamond panel).
@@
Moment closures approximating the third moment in terms of the first two moments inherently relies on two key assumptions. First, the third moment must be positive, as it represents triplet density and is calculated from the product of three non-negative densities [Eq. (2.6)]. Therefore, the closure approximation must also remain non-negative. The second assumption is the dynamical invariance under relabeling, implying that relabeling identical individuals should not alter the system dynamics [58]. Different closure methods balance these assumptions, with the power-2 closure we used in our computations offering an optimal trade-off between them. However, in certain parameter regimes, these assumptions may not hold, leading to notable discrepancies between IBM and SMD [59, 58]. Additionally, truncating the hierarchy of spatial moments at the second order may limit model accuracy for spatial structures with significant amounts of information in spatial moments at higher orders. Nevertheless, many previous studies successfully implemented moment closures to approximate spatial patterns and population dynamics in interacting particle system [60, 61, 32].
@@ 2.3 Field equation for the biomass density field
where we have omitted the time dependence on the right side. We now define the total number of plants in the population as $N=\int\langle n(\mathbf{x})\rangle\,\mathrm{d}\mathbf{x}=Z_{1}\times L^{2}$. Under the mean-field condition, the total number of plants in the population is denoted by $N_{\mathrm{h}}$. Notice that, similarly to what we obtained in the spatial moment dynamics equations in Section 2.2, Eq. (2.21) does not provide a closed description for the dynamics of the population density because it also includes its spatial covariance. We could write an equation for the dynamics of this spatial covariance and perform a moment closure approximation, as we did with the spatial moment dynamics equations. This approach is now much more complicated because the plant density is a field rather than a scalar quantity obtained from a spatial average. Therefore, we explore two alternative ways of treating these spatial correlations.
@@ 2.3.2 Beyond mean-field: introducing the effect of spatial covariances
All the quantities on the right side of Eq. (2.30) can be computed from IBM simulations. Alternatively, these quantities can also be obtained using the spatial moment dynamics (SMD) equations, which is computationally more efficient and makes it possible to have a closed dynamical equation for the vegetation density field without performing IBM simulations. Using the definitions of the spatial moments, we can rewrite (2.30) as,
@@ 3.1 Spatial patterns in the discrete model
Figure 1: IBM analysis using a Gaussian competition kernel. A) Snapshots of spatial patterns from IBM simulations. From left to right: segregated, uniform, and clustered patterns, denoted by diamond, circle, and triangle shapes, respectively. B) Total number of plants normalized by the homogenous (mean-field) expectation, $N_{\mathrm{h}}=(b-d)/g$. C) Radial correlation function at displacement zero $C(0)$ [36] as a function of the dispersal and competition ranges, $s$ and $r_{c}$ respectively. In both these heatmaps, we plot the steady-state values of $N/N_{\mathrm{h}}$ and $C(0)$, respectively, and depict time-averaged values computed up to $t=250$. D) Radial correlation function, $C(r)$, as a function of the separation distance, $r$, for three combinations of $s$ and $r_{c}$. These parameter combinations are indicated by the coordinates of the green symbols in the heat map and correspond to the respective patterns in A). The gray dashed line represents $C(r)=1$.
@@
Previous work has identified four possible types of spatial patterns depending on how these two scales compare to each other [36]. When both ranges are large, the dynamics are effectively non-spatial with a uniform spatial pattern, and the IBM converges to the classical logistic model. However, when the dispersal scale decreases, while keeping competition long-ranged, individuals tend to clump together forming a clustered spatial pattern. This scenario leads to smaller population sizes than in the non-spatial case. Likewise, when the competition range decreases while keeping dispersal long-range, individuals tend to overdisperse, forming a segregated pattern, which increases the total population size. We show the spatial snapshot of these different possible patterns in Fig. 1A. Finally, when both competition and dispersal are short-range, newborns cannot escape from competition with their ancestors and the population goes extinct. We recover these results integrating the SMD approximation in Eqs. (2.9)-(2.11) with the moment closure in Eq. (2.12) (Fig. 1B, C). In none of the scenarios discussed above, however, the spatial patterns are periodic (see radial correlation functions in Fig. 1D). Therefore, any of the continuous approximations we derived for the plant biomass density will return a homogeneous solution.
@@
To investigate whether regular patterns exist in our model, we solved the SMD equations using a top-hat kernel as defined in Eq. (2.3) and compared these predictions with results from simulations of the discrete IBM. The IBM simulations show the emergence of regular patterns when the dispersal range is short and the competition range long (Fig. 2). This result is in good agreement with previous work using various nonlocal extensions of the Fisher KPP equation with top-hat competition kernels [46, 52]. We find that the SMD equations accurately predict the value of the radial correlation function at pairwise distance zero (top panels in Fig. 3B, C), which is a proxy for local crowding. The SMD equations, however, predict less accurately the total number of plants in the population. Specifically, when the competition range is large, and the dispersal range is short, the SMD equations underestimate the total population size resulting from the IBM simulations (Fig. 3B, C).
@@
Figure 2: A) Long-time pattern snapshot with dispersal range $s=0.15$ and competition range $r_{c}=2.2$ (inside the region of high aggregation (labeled with the diamond marker) in the $C(0)$ heatmaps in Fig. 3). B) Fourier transform of the pattern snapshot. C) Radial correlation function, $C(r)$, of the pattern snapshot.
@@
This discrepancy in the total population size comes from the inaccurate prediction of the full radial correlation function provided by the SMD equations. In particular, SMD does not fully capture the amplitude of the oscillations in the radial correlation functions measured from IBM simulations (rightmost panel in the bottom row of Fig. 3B, C). This amplitude, especially when the radial correlation function becomes less than unity, is critical to determine how larger the population size grows when regular patterns form. More smaller the values of $C(r)$ indicate that the regions between clusters are less populated, which favors the growth of more individuals within each aggregate [47].
Figure 3: IBM analysis using a top-hat competition kernel. A) Snapshots of the spatial patterns from IBM simulations. From left to right: segregated, uniform, and regular aggregated patterns, denoted by triangle, circle, and diamond shapes, respectively. B) Top panels: SMD predictions for the radial correlation function at zero displacement, $C(0)$, and the total number of individuals normalized by the homogenous expectation, $N/N_{\mathrm{h}}$. In both these heatmaps, we plot the steady-state values of $C(0)$ and $N/N_{\mathrm{h}}$, respectively, which we computed using time-averaged values up to $t=250$. The bottom three panels in B) show the radial correlation function, $C(r)$, as a function of the separation distance $r$ for three specific combinations of $s$ and $r_{c}$. These parameter combinations are indicated by the coordinates of the red symbols in the heat maps and correspond to the respective patterns in A). The gray dashed lines represent $C(0)=1$. C) Same as in B) but computed directly from stochastic IBM simulations instead of SMD predictions.